OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3)with a slope of -6. Answer y = -6x + 20 y = -6x - 15 y = -6x + 15 y = -6x - 20

OpenStudy (anonymous):

B

OpenStudy (hoblos):

y+3 = -6(x +2 y= -6x -15